# On Uncertainty Calibration for Equivariant Functions

### v2

@@ 3.4 Error Bounds for Equivariant Models in Classification and Regression Tasks
We have $\text{err}(h)\geq\int_{F}p(Gx)\mathbb{V}_{Gx}[f]$.
The following theorem generalizes this to the setting of equivariance:
@@
We now only assume equivariance on $h$, that is, $h(\rho_{X}(g)x)=\rho_{Y}(g)h(x)$ where $g\in G$, $\rho_{X}$ and $\rho_{Y}$ are group representations associated with $X$ and $Y$. We will denote $\rho_{X}(g)x$ and $\rho_{Y}(g)y$ by $gx$ and $gy$, leaving the representation implicit. Assume $Y=\mathbb{R}^{n}$. Let Id be the identity. Define a matrix $Q_{Gx}\in\mathbb{R}^{n\times n}$ and $q(gx)\in\mathbb{R}^{n\times n}$ so that $\int_{G}q(gx)dg$ = Id by
$Q_{Gx}$ $=$ $\int_{G}p(gx)\rho_{Y}(g)^{T}\rho_{Y}(g)\alpha(x,g)dg$
@@
The error of $h$ has lower bound $\text{err}(h)\geq\int_{F}\int_{G}p(gx)\lVert f(gx)-g\mathcal{E}_{G}[f,x]\rVert_{2}^{2}\alpha(x,g)dgdx$.
@@ 3.5 Evidential Regression
This work studies the effect of equivariance on the ability to separate model and data-centric uncertainties. We now describe evidential regression, a learning framework that allocates mass between mean and dispersion under a surrogate loss; the allocation is not identifiable and is sensitive to loss shaping and misspecification. We review the relevant theory and notation from Amini et al. (2020).
Given $\left(y_{1},\ldots,y_{n}\right)\sim\mathcal{N}(\mu,\sigma^{2})$, we may impose priors
@@
The prediction definition is clear in that it represents the expected center of the Student t’s distribution used to fit the data. Aleatoric uncertainty represents the uncertainty in $P(Y|X)$, whereas the epistemic uncertainty represents uncertainty in the predictive law (the model). Practically, aleatoric uncertainty can be understood as the irreducible uncertainty due to inherent stochasticity in the data. In contrast, epistemic uncertainty is reducible uncertainty due to finite data, features, or misspecification. It vanishes in the limit of perfect knowledge and correct model class.
@@ 4 Invariant Classification Calibration
In this section, we present our results on bounds of uncertainty calibration error for classification for on invariant functions. In particular, we will show how miscalibration can be understood by examining the fibers of each uncertainty estimate and then studying the error on each of those fibers. The error on these fibers is related to the symmetries of the model and the data, which establishes a relationship between symmetry and model calibration. In the main, our bounds demonstrate how badly a function can become miscalibrated due to incorrect and extrinsic invariance. In particular, the lower and upper bounds are tightened by the dissent on individual fibers. Additionally, this section provides experimental results that shows how incorrect invariance can affect model calibration experimentally in domains where the theory is less tractable.
@@ 4.1 Classification Problem
Consider a function $f:X\to Y$ where $Y$ is a finite set of labels. Let $q:X\to\mathbb{R}$ be a probability density on the domain $X$. We define a function space $\mathscr{H}=\{h:X\to Y\times[0,1]\}$. If $h(x)=(h_{Y},h_{P})$ then $h_{P}$ represents the confidence estimate associated with the predicted label $h_{Y}$. The goal is to find the function $h\in\mathscr{H}$ that fits the function $f$ and to properly predicts its own confidence by minimizing the expected calibration error (Equation˜1, and Equation 2 in Guo et al. (2017)). Following Wang et al. (2024), we assume that the class $\mathscr{H}$ is arbitrarily expressive except that it is constrained to be equivariant with respect to a group $G$. In the classification setting, we specifically assume $h$ to be $G$-invariant. While not all classification problems are $G$-invariant, this is the case most commonly considered in the literature. Let $r(p)$ be the probability density such that $\mathbb{P}(p_{1}\leq h_{P}(x)\leq p_{2})=\int_{p_{1}}^{p_{2}}r(p)dp$. Equivalently, $r$ is the push-forward of $q$ over $h_{P}$. The expected calibration error is nominally defined
$\text{ECE}(h)=\mathbb{E}_{h_{P}}\left[\biggm|\mathbb{P}\left(f=h_{Y}|h_{P}=p\right)-p\biggm|\right]$ (1)
as in Guo et al. (2017). Intuitively, if a model has confidence $p$, then it should be accurate with probability $p$. This metric penalizes the discrepancy between accuracy and confidence averaged over all of confidences weighted by the push-forward density $r$. The definition in Guo et al. (2017) abuses notation slightly, in that the probability of any event drawn from a continuous random variable has probability zero, i.e. $p(h_{P}=p)=0$ for all $p$. We can rectify this by defining ECE as
$\text{ECE}(h)=\lim_{\varepsilon\to 0}\mathbb{E}_{p\sim r(p)}\left[\biggm|\mathbb{P}\left(f=h_{Y}|p-\varepsilon<h_{P}<p+\varepsilon\right)-p\biggm|\right],$ (2)
@@
We abbreviate $\mathbb{P}(f(x)=h_{Y}(x)|h_{P}(x)=p)$ with $\mathrm{Acc}_{p}(h)$. [^1] Some works refer to $\mathrm{Acc}_{p}(h)$ as the calibration function, e.g. Vaicenavicius et al. (2019) This is the true accuracy of the model when the predicted confidence is $p$. Hence $h$ is well calibrated at confidence $p$ when $\mathrm{Acc}_{p}(h)=p$, under confident when $\mathrm{Acc}_{p}(h)>p$ and overconfident when $\mathrm{Acc}_{p}(h)<p$. For the purposes of approximation, $|2\varepsilon|$ can be viewed as the bin width.
We briefly comment on the well-definedness of Equation˜2. First, $\mathbb{P}\left(f=h_{Y}|p-\varepsilon<h_{P}<p+\varepsilon\right)$ is well defined when $r(p)\neq 0$ for all $p\in[0,1]$. Moreover, we note that in general, if $\mathcal{C}=\{C=c\}$, it is not always permissible to define $\mathbb{P}(A|\mathcal{C})=\underset{\varepsilon\to 0}{\lim}\mathbb{P}(A|c-\varepsilon<C<c+\varepsilon)$. This is because we face contradictions when $\mathcal{C}=\{C=c\}=\{D=d\}$, but the random variables $C$ and $D$ have different densities defined with respect to different measures. This results in contradictions where $\mathbb{P}(A|\mathcal{C})=\underset{\varepsilon\to 0}{\lim}P(A|c-\varepsilon<C<c+\varepsilon)$ and $\mathbb{P}(A|\mathcal{C})=\underset{\varepsilon\to 0}{\lim}\mathbb{P}(A|d-\varepsilon<D<d+\varepsilon)$ but $\underset{\varepsilon\to 0}{\lim}\mathbb{P}(A|c-\varepsilon<C<c+\varepsilon)\neq\underset{\varepsilon\to 0}{\lim}\mathbb{P}(A|d-\varepsilon<D<d+\varepsilon)$, see for example the Borel-Kolmogorov Paradox (Kolmogorov, 1933; Jaynes, 2003) [^2] This paradox is most easily exemplified with the Great Circle Puzzle.. In other words, the probability density conditioned on an event with zero probability can only be specified with respect to a given reference measure that determines the probability density function being conditioned on. Therefore, we must specify a measure on $X$ so that the random variable $h_{p}$ has push-forward density $r(p)$ defined with respect to the push-forward measure. We express this with the following assumption.
@@
˜1 is sufficient for Equation˜2 to be uniquely defined. Our construction is supported by various disintegration theorems in the literature (e.g. Pachl, 1978; Chang & Pollard, 1997). For further background, see also Rokhlin (1949); Bogachev (2007). We also note that we do not need these well-definedness properties to hold in the special case where $r(p)$ is discrete or when we are computing approximations that treat $h_{P}$ as discrete. In each case, we average over the confidences (or confidence bins) with non-zero probability.
@@ 4.2 ECE Upper Bounds
The upper bound presented in the Proposition˜1 is loose without any further assumptions. We now show that assuming the function space is invariant allows us to tighten the upper bound. The proof uses iterated integration over the orbits and the set of orbit representatives (the fundamental domain). We review the integration formalism in Appendix˜A. Let us first make some assumptions in order for iterated integration on subsets of $X$ to be well defined.
@@
For a group $G$ acting on a domain $X$, the union of all pairwise intersections $\cup_{g_{1}\neq g_{2}}(g_{1}F\cap g_{2}F)$ have measure 0 and that $F$ and $Gx$ are differentiable manifolds for all $x\in X$. This holds for fibers $\mathcal{F}_{p}=h^{-1}(p)\subseteq X$ on which $G$ also acts. That is, if $F_{p}$ is a fundamental domain for the action of $G$ on $\mathcal{F}_{p}$, the union of all pairwise intersections $\cup_{g_{1}\neq g_{2}}(g_{1}F_{p}\cap g_{2}F_{p})$ have measure 0 and $F_{p}$ and $Gx$ are differentiable manifolds for all $x\in\mathcal{F}_{p}$.
@@
Denote the fiber $\mathcal{F}_{p}=h_{P}^{-1}(p)$. Denote the total density on a fiber $\mathcal{F}_{p}$ by $q(\mathcal{F}_{p})=\int_{\mathcal{F}_{p}}q(x)dx$ and the renormalized density by $q_{p}(x)=q(x)/q(\mathcal{F}_{p})$. Let $k_{p}(Gx)$ be the total dissent of an orbit on $\mathcal{F}_{p}$ with the renormalized probability $q_{p}(x)$. Let $F_{p}$ be a fundamental domain for the action of $G$ on $\mathcal{F}_{p}$. Assume that $h$ is incorrectly invariant under $G$ on each fiber of $\mathcal{F}_{p}$. In other words, $h$ satisfies $h(gx)=h(x)$ for $g\in G$, $x\in\mathcal{F}_{p}$, but $f(x)\neq f(gx)$ for some $x$, $g$. Let $P_{1}=\{p:\mathrm{Acc}_{p}(h)\leq 1/2\}$ and $P_{2}=\{p:\mathrm{Acc}_{p}(h)\geq 1/2\}$. Let $Gx^{*}$ be the orbit with the smallest nonzero total dissent $k(Gx^{*})$, i.e. $x^{*}=\underset{x\in X}{\arg\min}~k(Gx)$. Then ECE is bounded above by
@@
Observe that the upper bound for ECE is determined by the upper bound of Equation˜3 and Equation˜4 and is related to the accuracy of $h$ by the last two integrals in Equation˜4. In particular, the model $h$ that maximizes ECE satisfies $\mathrm{Acc}_{p}(h)=0$ on $P_{1}$ and $\mathrm{Acc}_{p}(h)=1-\int_{F_{p}}k(Gx,p)dx$ on $P_{2}$. Substituting these values into Equation˜4 gives upper bound
@@
This bound is non-vacuous. That is, the upper bound is tighter than $1$ since it accounts for the error caused by incorrect invariance along the subset of fibers where accuracy is at least $50\%$. By considering the orbit with the lowest total dissent, we can compute an upper bound that is tighter than $1$ even without knowing the error lower bound on each fiber, i.e. $\int_{F_{p}}k_{p}(Gx)dx$, or the fibers themselves $\mathcal{F}_{p}$. In other words, if we consider all of the orbits with incorrect invariance, then the ECE upper bound is only as tight as the smallest orbit-wise error bound will allow. Comparing with Theorem˜4, we see that, under the invariance assumption, the upper bound decreases by $k(Gx^{*})\int_{P_{2}}r(p)dp$. In other words, the bound is tight when there is nontrivial dissent on high mass orbits within fibers where the accuracy is greater than or equal to $1/2$. Conversely, the bound becomes trivial for models with correct invariance, or if the partition $P_{1}$ is a much larger subset than $P_{2}$.
@@
By assumption of invariance on $h_{P}$, the fibers of $[0,1]$ contain entire orbits. The integrated total dissent $\int_{F_{p}}k_{p}(Gx)dx$ is defined on the collection of orbits where the confidence is always given by $h_{p}(x_{p})=p$, but the label $h_{Y}(x_{p})$ itself may very. This is possible because points $x_{p_{1}}$ and $x_{p_{2}}$ may belong to distinct orbits which map to different distinct labels $y_{1}$ and $y_{2}$ under $h_{Y}$ but map to the same confidence $p$ under $h_{P}$.

### v4

@@ 3.4 Error Bounds for Equivariant Models in Classification and Regression Tasks
We have $\text{err}_{\text{reg}}(h)\geq\int_{F}p(Gx)\mathbb{V}_{Gx}[f]$.
theorem˜3 generalizes this to the setting of equivariance. Before stating the theorem, we establish how to lift an integral over an orbit to an integral over the group. Details on necessary hypotheses for this to be well defined are given in appendix˜A. Denote the stabilizer by $G_{x}=\{g:gx=x\}$. Denote by $a_{x}:G/G_{x}\to Gx$ the identification of the orbit $Gx$ and coset space $G/G_{x}$ with respect to the stabilizer. We have
$\int_{Gx}f(z)dz=\int_{G}f(gx)\alpha(g,x)dg$
where
We may now state the equivariant regression lower bound:
@@
Assume $h$ is equivariant: that is, $h(\rho_{X}(g)x)=\rho_{Y}(g)h(x)$ where $g\in G$, $\rho_{X}$ and $\rho_{Y}$ are group representations. Denote $\rho_{X}(g)x$ and $\rho_{Y}(g)y$ by $gx$ and $gy$. Let $Y=\mathbb{R}^{n}$ and Id be the identity. Define the matrix $Q_{Gx}\in\mathbb{R}^{n\times n}$ and $q(gx)\in\mathbb{R}^{n\times n}$ so that $\int_{G}q(gx)dg$ = Id by
$Q_{Gx}$ $=$ $\int_{G}p(gx)\rho_{Y}(g)^{T}\rho_{Y}(g)\alpha(x,g)dg,$
@@
The error of $h$ has lower bound $\text{err}_{\text{reg}}(h)\geq\int_{F}\int_{G}p(gx)\lVert f(gx)-g\mathcal{E}_{G}[f,x]\rVert_{2}^{2}\alpha(x,g)dgdx$.
@@ 3.5 Evidential Regression
This work studies the effect of equivariance on the ability to separate model and data-centric uncertainties. We now describe evidential regression, a learning framework that allocates mass between mean and dispersion under a surrogate loss. That is, a loss function determines how much uncertainty should be allocated to the model and how much should be allocated to the true dispersion (e.g. variance) of the data. The allocation is not identifiable and is sensitive to loss shaping and misspecification. We review the relevant theory and notation from section 2.3 in Hüllermeier and Waegeman (2021) for defining different sources of uncertainty, then Amini et al. (2020) for describing evidential regression.
The following definitions make the notions of model and data-centric uncertainties concrete:
Aleatoric Uncertainty:
Aleatoric uncertainty refers to the irreducible part of the uncertainty. Given spaces $X$ and $Y$ and an instance $x_{q}\in X$, the aleatoric uncertainty is the spread in $p(y|x_{q})$.
Epistemic Uncertainty:
Model uncertainty and approximation uncertainty, on the other hand, are subsumed under the notion of epistemic uncertainty. Let the spaces $X$ and $Y$ be the same as before. Let $l:Y\times Y\to\mathbb{R}$ be the loss function and let $f^{*}$ be the associated point-wise Bayes predictor defined as
$f^{*}(x):=\underset{\hat{y}\in Y}{\arg\min}\int_{Y}l(y,\hat{y})dP(y|x).$ (1)
Epistemic uncertainty is the uncertainty due to the lack of knowledge of the perfect predictor equation˜14.
Evidential Regression.
We now describe evidential regression, which prescribes a specific parameterization of the two uncertainties. Given $\left(y_{1},\ldots,y_{n}\right)\sim\mathcal{N}(\mu,\sigma^{2})$, we may impose priors
@@
The prediction definition is clear in that it represents the expected mean of the Student’s t distribution used to fit the data. Aleatoric uncertainty represents the uncertainty in $P(y|x)$, whereas the epistemic uncertainty represents uncertainty in the predictive law (the model). Practically, aleatoric uncertainty can be understood as the irreducible uncertainty due to inherent stochasticity in the data. In contrast, epistemic uncertainty is reducible uncertainty due to finite data, features, or misspecification. It vanishes in the limit of perfect knowledge and correct model class.
@@ 4 Invariant Classification Calibration
In this section, we present our results on bounds of uncertainty calibration error for classification on invariant functions. In particular, we will show how miscalibration can be understood by examining the fibers of each uncertainty estimate and then studying the error on each of those fibers. The error on these fibers is related to the symmetries of the model and the data, which establishes a relationship between symmetry and model calibration. In the main, our bounds demonstrate how badly a function can become miscalibrated due to incorrect and extrinsic invariance. In particular, the lower and upper bounds are tightened by the dissent on individual fibers. Additionally, this section provides experimental results that show how incorrect invariance can affect model calibration in domains where the theory is less tractable. For practitioners who are more interested in empirical results and conclusions, experiments begin in section˜4.7.
@@ 4.1 Classification Problem
Consider a function $f:X\to Y$ where $Y$ is a finite set of labels. Let $q:X\to\mathbb{R}$ be a probability density on the domain $X$. We define a function space $\mathscr{H}=\{h:X\to Y\times[0,1]\}$. If $h(x)=(h_{Y},h_{P})$ then $h_{P}$ represents the confidence estimate associated with the predicted label $h_{Y}$. The goal is to find the function $h\in\mathscr{H}$ that fits the function $f$ and to properly predict its own confidence by minimizing the expected calibration error (equation˜2, and Equation 2 in Guo et al. (2017)). Following Wang et al. (2024), we assume that the class $\mathscr{H}$ is arbitrarily expressive except that it is constrained to be equivariant with respect to a group $G$. In the classification setting, we specifically assume $h$ to be $G$-invariant. While not all classification problems are $G$-invariant, this is the case most commonly considered in the literature. Let $r(p)$ be the probability density such that $\mathbb{P}(p_{1}\leq h_{P}(x)\leq p_{2})=\int_{p_{1}}^{p_{2}}r(p)dp$. Equivalently, $r$ is the push-forward of $q$ over $h_{P}$. The expected calibration error is nominally defined
$\text{ECE}(h)=\mathbb{E}_{h_{P}}\left[\biggm|\mathbb{P}\left(f=h_{Y}\mid h_{P}=p\right)-p\biggm|\right]$ (2)
as in Guo et al. (2017). Intuitively, if a model has confidence $p$, then it should be accurate with probability $p$. This metric penalizes the discrepancy between accuracy and confidence averaged over all of confidences weighted by the push-forward density $r$. The definition in Guo et al. (2017) abuses notation slightly, in that the probability of any event drawn from a continuous random variable has probability zero, i.e., $p(h_{P}=p)=0$ for all $p$. We can rectify this by defining ECE as
$\text{ECE}(h)=\lim_{\varepsilon\to 0}\mathbb{E}_{p\sim r(p)}\left[\biggm|\mathbb{P}\left(f=h_{Y}\mid p-\varepsilon<h_{P}<p+\varepsilon\right)-p\biggm|\right],$ (3)
@@
We abbreviate $\mathbb{P}(f(x)=h_{Y}(x)|h_{P}(x)=p)$ as $\mathrm{Acc}_{p}(h)$. [^1] Some works refer to $\mathrm{Acc}_{p}(h)$ as the calibration function, e.g. Vaicenavicius et al. (2019) This is the true accuracy of the model when the predicted confidence is $p$. Hence $h$ is well calibrated at confidence $p$ when $\mathrm{Acc}_{p}(h)=p$, under confident when $\mathrm{Acc}_{p}(h)>p$ and overconfident when $\mathrm{Acc}_{p}(h)<p$. For the purposes of approximation, $|2\varepsilon|$ can be viewed as the bin width.
We briefly comment on the well-definedness of equation˜3. First, $\mathbb{P}\left(f=h_{Y}|p-\varepsilon<h_{P}<p+\varepsilon\right)$ is well defined when $r(p)\neq 0$ for all $p\in[0,1]$. Moreover, we note that in general, if $\mathcal{C}=\{C=c\}$, it is not always permissible to define $\mathbb{P}(A|\mathcal{C})=\underset{\varepsilon\to 0}{\lim}\mathbb{P}(A|c-\varepsilon<C<c+\varepsilon)$. This is because we face contradictions when $\mathcal{C}=\{C=c\}=\{D=d\}$, but the random variables $C$ and $D$ have different densities defined with respect to different measures. This results in contradictions where $\mathbb{P}(A|\mathcal{C})=\underset{\varepsilon\to 0}{\lim}P(A|c-\varepsilon<C<c+\varepsilon)$ and $\mathbb{P}(A|\mathcal{C})=\underset{\varepsilon\to 0}{\lim}\mathbb{P}(A|d-\varepsilon<D<d+\varepsilon)$ but $\underset{\varepsilon\to 0}{\lim}\mathbb{P}(A|c-\varepsilon<C<c+\varepsilon)\neq\underset{\varepsilon\to 0}{\lim}\mathbb{P}(A|d-\varepsilon<D<d+\varepsilon)$, see for example the Borel-Kolmogorov Paradox (Kolmogorov, 1933; Jaynes, 2003) [^2] This paradox is most easily exemplified with the Great Circle Puzzle.. In other words, the probability density conditioned on an event with zero probability can only be specified with respect to a given reference measure that determines the probability density function being conditioned on. Therefore, we must specify a measure on $X$ so that the random variable $h_{p}$ has push-forward density $r(p)$ defined with respect to the push-forward measure. We express this with the following assumption.
@@
˜1 is sufficient for equation˜3 to be uniquely defined. Our construction is supported by various disintegration theorems in the literature (e.g. Pachl, 1978; Chang and Pollard, 1997). For further background, see also Rokhlin (1949); Bogachev (2007). We also note that we do not need these well-definedness properties to hold in the special case where $r(p)$ is discrete or when we are computing approximations that treat $h_{P}$ as discrete. In each case, we average over the confidences (or confidence bins) with non-zero probability.
@@ 4.2 ECE Upper Bounds
The upper bound presented in the proposition˜1 is loose without any further assumptions. We now show that assuming the function space is invariant allows us to tighten the upper bound. The proof uses iterated integration over the orbits and the set of orbit representatives (the fundamental domain). Let us first make some assumptions in order for iterated integration on subsets of $X$ to be well defined.
@@
For a group $G$ acting on a domain $X$, the union of all pairwise intersections $\cup_{g_{1}\neq g_{2}}(g_{1}F\cap g_{2}F)$ have measure 0 and that $F$ and $Gx$ are differentiable manifolds for all $x\in X$. This holds for fibers $\mathcal{F}_{p}=h^{-1}(p)\subseteq X$ on which $G$ also acts. That is, if $F_{p}$ is a fundamental domain for the action of $G$ on $\mathcal{F}_{p}$, the union of all pairwise intersections $\cup_{g_{1}\neq g_{2}}(g_{1}F_{p}\cap g_{2}F_{p})$ have measure 0 and $F_{p}$ and $Gx$ are differentiable manifolds, eventually with boundary, for all $x\in\mathcal{F}_{p}$. Note that example˜1 satisfies all the above.
@@
Denote the fiber $\mathcal{F}_{p}=h_{P}^{-1}(p)$. Denote the total density on a fiber $\mathcal{F}_{p}$ by $q(\mathcal{F}_{p})=\int_{\mathcal{F}_{p}}q(x)dx$ and the renormalized density by $q_{p}(x)=q(x)/q(\mathcal{F}_{p})$. Let $k_{p}(Gx)$ be the total dissent of an orbit on $\mathcal{F}_{p}$ with the renormalized probability $q_{p}(x)$. Let $F_{p}$ be a fundamental domain for the action of $G$ on $\mathcal{F}_{p}$. Assume that $h$ is incorrectly invariant under $G$ on each fiber of $\mathcal{F}_{p}$. In other words, $h$ satisfies $h(gx)=h(x)$ for $g\in G$, $x\in\mathcal{F}_{p}$, but $f(x)\neq f(gx)$ for some $x$, $g$. Let $P_{1}=\{p\colon\mathrm{Acc}_{p}(h)\leq 1/2\}$ and $P_{2}=\{p\colon\mathrm{Acc}_{p}(h)\geq 1/2\}$. Let $Gx^{*}$ be the orbit with the smallest nonzero total dissent $k(Gx^{*})$, i.e., $x^{*}=\underset{x\in X}{\arg\min}~k(Gx)$. Then ECE is bounded above by
@@
Observe that the upper bound for ECE is determined by the upper bound of equation˜4 and equation˜5 and is related to the accuracy of $h$ by the last two integrals in equation˜5. In particular, the model $h$ that maximizes ECE satisfies $\mathrm{Acc}_{p}(h)=0$ on $P_{1}$ and $\mathrm{Acc}_{p}(h)=1-\int_{F_{p}}k(Gx,p)dx$ on $P_{2}$. Substituting these values into equation˜5 gives upper bound
@@
This bound is non-vacuous. That is, the upper bound is tighter than $1$ since it accounts for the error caused by incorrect invariance along the subset of fibers where accuracy is at least $50\%$. By considering the orbit with the lowest total dissent, we can compute an upper bound that is tighter than $1$ even without knowing the error lower bound on each fiber, i.e., $\int_{F_{p}}k_{p}(Gx)dx$, or the fibers themselves $\mathcal{F}_{p}$. In other words, if we consider all of the orbits with incorrect invariance, then the ECE upper bound is only as tight as the smallest orbit-wise error bound will allow. Comparing with theorem˜4, we see that, under the invariance assumption, the upper bound decreases by $k(Gx^{*})\int_{P_{2}}r(p)dp$. In other words, the bound is tight when there is nontrivial dissent on high mass orbits within fibers where the accuracy is greater than or equal to $1/2$. Conversely, the bound becomes trivial for models with correct invariance, or if the partition $P_{1}$ is a much larger subset than $P_{2}$.
@@
By assumption of invariance on $h_{P}$, the fibers of $[0,1]$ contain entire orbits. The integrated total dissent $\int_{F_{p}}k_{p}(Gx)dx$ is defined on the collection of orbits where the confidence is always given by $h_{p}(x_{p})=p$, but the label $h_{Y}(x_{p})$ itself may vary. This is possible because points $x_{p_{1}}$ and $x_{p_{2}}$ may belong to distinct orbits which map to different distinct labels $y_{1}$ and $y_{2}$ under $h_{Y}$ but map to the same confidence $p$ under $h_{P}$.